Jonathan is 15 years older than Jean. Six years ago, Jonathan was six times as old as Jean. What is Jonathan present age?
Answer by satyareddy22(84)

Let the age of Jonathan be x and age of Jean be y.
x-y=15-----(1)
Six years ago,
x-6=6(y-6)
x-6=6y-36=>x-6y=-36=6=>x-6y=-30-----(2)
solving (1)&(2)
5y=45=>y=45/5=9=>y=9yrs
x=15=9=24=>x=24yrs
Jonathan present age=24years